OpenStudy (anonymous):

Calculate the energy needed to change 200.0 g of ice from -15.0° C to water at 10.0° C.

OpenStudy (anonymous):

Besides heat required for increase in temperature of ice(water) ,we also need heat for changing state from ice at 0 degrees to water at 0 degrees. So heat required = ΔT*s*m + mL ΔT=25 m=0.2kg s=specific heat capacity of water (or ice) = 4186 J/Kg/K L=Latent heat of water =334000 J/kg

OpenStudy (anonymous):

So would my equation be.. \[(25 \times 4186 \times 0.2) + (0.2 \times 334000)\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So my answer would be 87,730?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thank you